# Accurately simulating core-collapse self-interacting dark matter halos

### v2

@@ 3.1 Relation between softening length and variable time steps
Figure 1: Interplay of softening length and adaptive time steps for energy conservation. Similar to Fig. 17, we show the time evolution of the mass enclosed within 10 pc (upper panel) and the energy conservation (lower panel) for simulations of collisionless DM. The shaded regions indicate the uncertainties estimated based on shot noise. The simulations employ an adaptive time step, except for the results given by the yellow curve. More details can be found in Tab. 1, the corresponding names are: A, B, C, D, E, and F.
Here, we investigate the size of the gravitational softening length and its relation to the use of an adaptive time-stepping scheme. For this purpose, we consider CDM simulations only.
From Fig. 1, we can see that a too small gravitational softening length can lead to a loss of energy over the course of the simulation (compare light green and light blue). However, when fixing the adaptive time steps to a constant value for all particles over the whole simulation (yellow curve), the error in energy conservation vanishes. This can be understood by a too small softening length causing more frequent changes of time steps, implying a larger error in total energy. The energy error related to the time steps is a consequence of choosing the time steps in a non-time-symmetric way (Dehnen, 2017).
We note that we have chosen a fairly small time step for the simulation with the fixed time step. This makes sure that all particles are evolved on a sufficiently small time step, but at the same time reduces integration errors that are not directly related to the change of time steps as well. If we would instead choose a much larger constant time step for all particles, this leads to an increase in total energy by amplifying the errors that arise from the asymmetric tree evaluation that we do in OpenGadget3, implying that it decreases when shrinking the opening angle for the tree nodes (see also the explanation given by Fischer et al., 2024a, at the end of their Sect. 3.2.1). In addition, we have also checked that the energy conservation error for the simulation with the small softening length and the variable time step (light blue) does not arise from the error related to the asymmetric tree evaluation, i.e. we have explicitly checked that decreasing the opening angle does not affect the error in energy conservation. In conclusion, the statement from the previous paragraph about the energy conservation error arising from the asymmetric change of time steps is valid.
@@ 3.2 Sampling radius of ICs
Figure 2: Variation of the maximum sampling radius for the ICs. Analogously to Fig. 17, the mass enclosed within 10 pc (upper panel) and the energy conservation (lower panel) are shown as a function of time. The results for ICs sampled up to $10\,r_{\mathrm{s}}$ (orange and light blue) and $15\,r_{\mathrm{s}}$ (purple and dark blue) are shown. We note that the lower resolution simulations (orange and purple) employ a softening length of $\epsilon=6.4$ pc, whereas the higher resolution ones (light and dark blue) employ a softening length of $\epsilon=1.0$ pc. All parameters for the shown simulations, G, J, M, and O are given in Tab. 1.
As a next step, we investigate the role of the ICs. In particular, an approximation that is commonly made when sampling a halo model like the NFW profile.
Sampling the ICs for a density profile that does imply an infinite mass, such as the NFW profile, requires truncating the halo at a radius $r_{\mathrm{cut}}$. The truncation of the halo can influence its gravothermal evolution and, for example, speed up the collapse for satellite galaxies, where the outer regions of the halo are removed due to tidal force (e.g. Nishikawa et al., 2020; Sameie et al., 2020). In contrast to this physical truncation, we investigate the role of the artificially set truncation radius for SIDM simulations.
In Fig. 2, we test the influence of $r_{\mathrm{cut}}$ on the evolution of the halo. A comparison of simulations with ICs sampled up to $10\,r_{\mathrm{s}}$ and $15\,r_{\mathrm{s}}$ reveals a significant deviation in the evolution. While the difference for the mass within 10 pc is fairly large towards the end of our simulations, this can also be interpreted as a much smaller deviation in collapse time. Depending on how accurately the evolution should be predicted, $r_{\mathrm{cut}}=10\,r_{\mathrm{s}}$ could be too small, and a larger value would be preferable.
@@ 3.3 The role of resolution and softening length
Figure 3: Variation of resolution and gravitational softening length. Similar to Fig. 17, we show the enclosed mass within 10 pc and the energy conservation as a function of time. The simulations J, N, O, and Q as given in Tab. 1 are shown.
We now study the role of the numerical resolution, i.e. the number of numerical particles that are used to resolve the DM halo. At the same time, we consider two different choices for the gravitational softening length.
The results for four of our SIDM simulations are given in Fig. 3. From a comparison of the purple and black curves, we can see that the collapse time is interestingly increasing slightly for a higher resolution. An effect that could contribute to this is gravitational two-body relaxation being stronger in simulations with lower resolution. But also an increased effective heat conduction due to a large SIDM kernel size compared to the mean free path can contribute to this, as we discuss in Sect. 3.5. We note that the difference between those two runs may not be caused by inaccurate energy conservation. In fact, the simulation indicated by the purple curve is slightly gaining energy, which would slow down the evolution (see also Sect. 3.4). However, this is in contrast to what we find when we compare it to the black curve.
Moreover, we do not find clear evidence that the gravitational softening length has a significant impact on the evolution for the values that we have tested. This suggests that the results are not very sensitive to the gravitational softening as long as the softening length is chosen within a reasonable range. The very slight difference between the green and blue curves for the enclosed mass might be related to the differences in energy conservation.
A common choice is to set the gravitational softening length according to Power et al. (2003) or van den Bosch & Ogiya (2018). However, the optimal gravitational softening length for an SIDM halo in the collapse phase might be smaller than the one for the pre-collapsing halo, e.g. an NFW halo. This is especially the case if one is interested in the inner regions of the collapsing object, i.e. fairly small length scales. As a consequence, one may eventually want to choose the softening length somewhat smaller, closer to under-softening than recommended according to criteria based on CDM simulations as a compromise.
@@ 3.4 The role of energy conservation
Figure 4: The impact of energy conservation on the evolution time of the halo. Following Fig. 17, we show the enclosed mass within 10 pc (upper panel) and the energy conservation (lower panel) as a function of time. The shown simulations differ in softening length, causing differences in the accuracy of energy conservation, as can be understood with the explanations given in Sect. 3.1. All parameters for the displayed simulations J and L are given in Tab. 1.
@@
In Fig. 4, we show two simulations that differ in energy conservation. The one with the smaller softening length (red) loses energy in line with the explanation given in Sect. 3.1. It loses roughly 0.6% of its total energy, resulting in a speed-up of the evolution of about 4.2%, when compared to the simulation with the larger softening length (purple). We note that the larger softening length is not directly responsible for a significant slowdown of the collapse for the corresponding simulation (purple), as we have found in Sect. 3.3. Overall, this illustrates that an accurate prediction of the collapse time requires a fairly accurate conservation of total energy.
@@ 3.5 SIDM kernel size
Figure 5: The evolution of an isolated halo as a function of time for two different choices of $N_{\mathrm{ngb}}$. The upper panel gives the mass enclosed within $10\,\mathrm{pc}$ and the lower panel displays the energy conservation as a function of time. The choice of $N_{\mathrm{ngb}}=384$ corresponds to twice the kernel size compared to the SIDM computations for $N_{\mathrm{ngb}}=48$. Table 1 gives the parameters employed for the displayed simulations J and K.
Figure 6: SIDM kernel size divided by the mean free path as a function of radius for several times. The results, computed according to Eq. (2), are shown for several simulations with different resolutions. All simulations were run with $N_{\mathrm{ngb}}=48$. More details for the shown simulations J, O, and Q are given in Tab. 1.
It may have been Koda & Shapiro (2011) who first pointed out that SIDM $N$-body simulations should resolve the mean free path set by the local density and cross-section. Fischer & Sagunski (2024) found for a set-up rather different from a halo simulation that the ratio of the SIDM kernel size, $h$, to the mean free path, $l$, actually matters. Here, we want to investigate the relevance of this for simulating the collapse phase of SIDM halos.
For this purpose, we rerun one of our simulations, increasing the neighbour number from $N_{\mathrm{ngb}}=48$ to $N_{\mathrm{ngb}}=384$. This implies an increase in the kernel size by a factor of two. In Fig. 5, we show the results for the mass enclosed within 10 pc and the energy conservation. For almost all of the evolution, there is basically no difference between the two simulations. Only at very late stages, where $h/l$ has increased a lot at small radii, due to the collapse, we find that the simulation with the larger value for $N_{\mathrm{ngb}}$ reaches high densities earlier. This could be explained by a larger SIDM kernel size leading to an artificially strong effective heat conduction.
We compute the ratio between kernel size, $h$, and mean free path, $l$, following Fischer et al. (2024a)
@@
Here, $\sigma_{\mathrm{eff}}$ is defined as the effective cross-section given by Eq. 11. In the case of a velocity-independent isotropic cross-section, $\sigma_{\mathrm{eff}}=\sigma$. The result is given for various simulations and times as a function of radius in Fig. 6. We find that a kernel size that is even a few multiples larger than the mean free path appears to be unproblematic (compare also Fig. 3). In contrast, ratios exceeding an order of magnitude appear to be problematic. However, in general, it depends on the simulation set-up how problematic a specific value for $h/l$ is; in particular, it should depend on the velocity dispersion gradient. In this sense, stages of the gravothermal collapse later than those that we have simulated could be even more problematic. Furthermore, we want to add that there might be a mild dependence on how the geometric factor $\Lambda$ for the interaction probability and drag force is computed (see Eqs. 9, 13 and B4 by Fischer et al., 2021a), e.g. the chosen kernel function may matter.
Finally, we want to point out that the error arising from a too large value for $h/l$ does not show up in the energy conservation of the simulation. Hence, this is a good example of where energy conservation fails as a measure of the simulation’s accuracy.
@@ 3.6 Using a minimal time step
Figure 7: Employing a minimum time-step. As in Fig. 1, we show the enclosed mass within 10 pc and the energy conservation for various simulations. This time, we investigate how much the simulations are affected when employing a minimum time step. The brighter lines indicate simulations without a minimum time-step, i.e. the time-step can decrease further. In contrast the darker lines feature a minimum time-step of $\Delta t_{\mathrm{min}}=2.8\times 10^{-5}\,\mathrm{Gyr}$. The displayed simulations J, N, O, Jt, Nt, and Ot are described in Tab. 1.
@@
Figure 8: Density, velocity dispersion and Knudsen number for simulations employing a minimum time step. The purple lines give the results for the simulation with $N=2\times 10^{6}$ and $\epsilon=6.4\,\mathrm{pc}$ (grey) and $\Delta t_{\mathrm{min}}=2.8\times 10^{-5}\,\mathrm{Gyr}$ and the blue lines are for the simulation with $N=1\times 10^{7}$ and $\epsilon=1.0\,\mathrm{pc}$ and $\Delta t_{\mathrm{min}}=2.8\times 10^{-5}\,\mathrm{Gyr}$. These are the same simulations as in Fig. 7. The top panel gives the density as a function of radius for different times. We note that these times are during the collapse phase, and for most of them, the energy conservation error is sizeable. The middle panel gives the velocity dispersion as a function of radius, and the bottom panel displays the Knudsen number (Eq. (3)). All parameters for the shown simulations Jt and Ot are given in Tab. 1.
@@
Next, we study how imposing a minimum time step impacts the simulation results in the collapse phase. The time step criterion for SIDM, but also gravity, requires smaller and smaller time steps while the halo is collapsing. In particular, the SIDM time step criterion becomes prohibitively small. This poses a challenge for cosmological simulations, as one aims to run until a specific redshift, e.g. $z=0$. A strategy to complete such a simulation, although some halos are collapsing, is to set a minimum allowed time step despite the time step criteria, i.e. evolving some particles on too large time steps.
@@
Following this idea, we set a minimum time step for the late evolution phase. In practice, we continue some of our simulations with such a minimum time step. In our case we choose $\Delta t_{\mathrm{min}}=2.8\times 10^{-5}\,\mathrm{Gyr}$. The results for the simulations for which we test this are displayed in Fig. 7. Here, we show the mass enclosed within 10 pc and the energy conservation error. For the late collapse phase, we find that the simulations with the minimum time step agree well with the ones that can contain particles on even smaller time steps. Having a minimum time step in place, we can continue the simulations to even later stages with higher central densities. However, at some point, the central density stops growing while the total energy is no longer conserved but increases drastically. Interestingly, the increase in total energy seems to be very sensitive to the gravitational softening length.
@@
Studying the mass enclosed within 10 pc only may give an incomplete picture of the impact of imposing a minimum time step. Instead, we consider the density, the velocity dispersion and the Knudsen number,
$Kn=\sqrt{\frac{4\pi}{\rho\,\nu^{2}}}\left(\frac{\sigma_{\mathrm{eff}}}{m}\right)^{-1}\,.$ (3)
@@
We use the effective cross-section, $\sigma_{\mathrm{eff}}$, it is given by Eq. 11. Moreover, for a velocity-independent isotropic cross-section, $\sigma_{\mathrm{eff}}=\sigma$. We compute those quantities as a function of radius and show them for two of our continued simulations in Fig. 8. For the first time that we show ($4.7\,\mathrm{Gyr}$), the two simulations give fairly similar results. However, at later times they deviate from each other, sometimes leading to huge differences, e.g. for the velocity dispersion. This demonstrates that measuring quantities from simulations of collapsing halos where a minimum time step had been imposed can be far off.
@@
Nevertheless, the simulation Jt shown in purple in Fig. 8 might be more accurate as the energy conservation is much better. The results for the very late collapse phase, where we find an increase of the central density compared to maximum core formation by a factor of $6\times 10^{5}$, also appear to be qualitatively in agreement with the findings in Fig. 3 by Balberg et al. (2002) for the gravothermal fluid model.
@@
In Sect. 3.7, we will comment more on the errors arising from imposing a minimum time step. In detail, we will look at the enclosed mass in projection, which is relevant for gravitational lensing.

### v4

@@ 3.1 Relation between softening length and variable time steps
Figure 1: Interplay of softening length and adaptive time steps for energy conservation. We show the time evolution of the mass enclosed within 10 pc (upper panel) and the energy conservation (lower panel) for simulations of collisionless DM. The shaded regions indicate the uncertainties estimated based on shot noise. The simulations employ an adaptive time step, except for the results given by the yellow curve. More details can be found in Table 1; the corresponding names are A, B, C, D, E, and F.
Here we investigate the size of the gravitational softening length and its relation to the use of an adaptive time-stepping scheme. For this purpose, we consider only CDM simulations.
In Fig. 1 we can see that a too small gravitational softening length can lead to a loss of energy over the course of the simulation (compare light green and light blue). However, when fixing the adaptive time steps to a constant value for all particles over the whole simulation (yellow curve), the error in energy conservation vanishes. This can be understood by a too small softening length, which causes more frequent changes in time steps, thus implying a larger error in total energy. The energy error related to the time steps is a consequence of choosing the time steps in a non-time-symmetric way (Dehnen, 2017).
We note that we have chosen a fairly small time step for the simulation with the fixed time step. This ensures that all particles are evolved on a sufficiently small time step, but at the same time reduces integration errors that are not directly related to the change in time steps as well. If we instead choose a much larger constant time step for all particles, this leads to an increase in total energy by amplifying the errors that arise from the asymmetric tree evaluation that we do in OpenGadget3, implying that it decreases when shrinking the opening angle for the tree nodes (see also the explanation given by Fischer et al., 2024a, at the end of their Sect. 3.2.1). In addition, we also checked that the energy conservation error for the simulation with the small softening length and the variable time step (light blue) does not arise from the error related to the asymmetric tree evaluation, i.e. we explicitly checked that decreasing the opening angle does not affect the error in energy conservation. In conclusion, the statement from the previous paragraph about the energy conservation error arising from the asymmetric change of time steps is valid.
@@ 3.2 Sampling radius of ICs
Figure 2: Variation of the maximum sampling radius for the ICs. Analogously to Fig. 1, the mass enclosed within 10 pc (upper panel) and the energy conservation (lower panel) are shown as a function of time. The results for ICs sampled up to $10\,r_{\mathrm{s}}$ (orange and light blue) and $15\,r_{\mathrm{s}}$ (purple and dark blue) are shown. We note that the lower resolution simulations (orange and purple) employ a softening length of $\epsilon=6.4$ pc, whereas the higher resolution simulations (light and dark blue) employ a softening length of $\epsilon=1.0$ pc. All parameters for the shown simulations (G, J, M, and O) are given in Table 1.
As a next step, we investigated the role of the ICs. In particular, we made an approximation that is commonly made when sampling a halo model such as the NFW profile.
Sampling the ICs for a density profile that implies an infinite mass, such as the NFW profile, requires truncating the halo at a radius $r_{\mathrm{cut}}$. The truncation of the halo can influence its gravothermal evolution and, for example, can speed up the collapse for satellite galaxies, where the outer regions of the halo are removed due to tidal force (e.g. Nishikawa et al., 2020; Sameie et al., 2020). In contrast to this physical truncation, we investigated the role of the artificially set truncation radius for SIDM simulations.
In Fig. 2 we test the influence of $r_{\mathrm{cut}}$ on the evolution of the halo. A comparison of simulations with ICs sampled up to $10\,r_{\mathrm{s}}$ and $15\,r_{\mathrm{s}}$ reveals a significant deviation in the evolution. Although the simulations agree with each other during the core formation and early collapse phase within the estimated Poisson noise based on the number of simulation particles within 10 pc, they start to deviate at later times. Here the enclosed mass grows super-exponentially, and the relative uncertainty on the inferred mass goes down. At the same time, the difference between the simulations in terms of the enclosed mass becomes relatively large, although the deviation in the collapse time is small. Because of the super-exponential nature of the enclosed mass growth in the deep core-collapse phase, a small deviation in the collapse time results in a large difference in the mass. Depending on how accurately the evolution should be predicted, $r_{\mathrm{cut}}=10\,r_{\mathrm{s}}$ could be too small, and a higher value would be preferable.
@@ 3.3 The role of resolution and softening length
Figure 3: Variation of resolution and gravitational softening length. Similar to Fig. 1, we show the enclosed mass within 10 pc and the energy conservation as a function of time. The simulations J, N, O, and Q as given in Table 1 are shown.
We now look at the role of the numerical resolution, i.e. the number of numerical particles that are used to resolve the DM halo. At the same time, we consider two different choices for the gravitational softening length.
The results for four of our SIDM simulations are given in Fig. 3. Interestingly, from a comparison of the purple and black curves, we can see that the collapse time is increasing slightly with higher resolution. An effect that could contribute to this is gravitational two-body relaxation being stronger in simulations with lower resolution. An increased effective heat conduction due to a large SIDM kernel size compared to the mean free path can also contribute to this, as we discuss in Sect. 3.5. We note that the difference between these two runs may not be caused by inaccurate energy conservation. In fact, the simulation indicated by the purple curve is slightly gaining energy, which would slow down the evolution (see also Sect. 3.4). However, this is in contrast to what we find when we compare it with the black curve.
Moreover, we do not find clear evidence that the gravitational softening length has a significant impact on the evolution for the values that we have tested. This suggests that the results are not very sensitive to gravitational softening as long as the softening length is chosen within a reasonable range. The very slight difference between the green and blue curves for the enclosed mass might be related to differences in energy conservation.
A common choice is to set the gravitational softening length according to Power et al. (2003) or van den Bosch & Ogiya (2018). However, the optimal gravitational softening length for an SIDM halo in the collapse phase might be smaller than that for the pre-collapsing halo (e.g. an NFW halo). This is especially the case if one is interested in the inner regions of the collapsing object, i.e. fairly small length scales. As a consequence, as a compromise one may eventually want to choose a softening length that is somewhat smaller, closer to undersoftening than recommended according to the criteria based on CDM simulations.
@@ 3.4 The role of energy conservation
Figure 4: Impact of energy conservation on the evolution time of the halo. Following Fig. 1, we show the enclosed mass within 10 pc (upper panel) and the energy conservation (lower panel) as a function of time. The shown simulations differ in softening length, causing differences in the accuracy of energy conservation (see explanations in Sect. 3.1). All parameters for the displayed simulations J and L are given in Table 1.
@@
In Fig. 4 we show two simulations that differ in energy conservation. The one with the smaller softening length (red) loses energy in line with the explanation given in Sect. 3.1. It loses approximately 0.6% of its total energy, resulting in a speed-up of the evolution of about 4.2%, when compared to the simulation with the larger softening length (purple). We note that the larger softening length is not directly responsible for a significant slowdown of the collapse for the corresponding simulation (purple), as we found in Sect. 3.3. Overall, this illustrates that an accurate prediction of the collapse time requires a fairly accurate conservation of total energy.
@@ 3.5 SIDM kernel size
Figure 5: Evolution of an isolated halo as a function of time for two different choices of $N_{\mathrm{ngb}}$. The upper panel gives the mass enclosed within $10\,\mathrm{pc}$ and the lower panel displays the energy conservation as a function of time. The choice of $N_{\mathrm{ngb}}=384$ corresponds to twice the kernel size compared to the SIDM computations for $N_{\mathrm{ngb}}=48$. Table 1 gives the parameters employed for the displayed simulations J and K.
Figure 6: SIDM kernel size divided by the mean free path as a function of radius for several times. The results, computed according to Eq. (2), are shown for several simulations with different resolutions. All simulations were run with $N_{\mathrm{ngb}}=48$. More details for the shown simulations J, O, and Q are given in Table 1.
It may have been Koda & Shapiro (2011) who first pointed out that SIDM $N$-body simulations should resolve the mean free path set by the local density and cross-section. Fischer & Sagunski (2024) found for a set-up rather different from a halo simulation that the ratio of the SIDM kernel size, $h$, to the mean free path, $l$, actually matters. Here we investigate the relevance of this for simulating the collapse phase of SIDM halos.
For this purpose, we reran one of our simulations, increasing the neighbour number from $N_{\mathrm{ngb}}=48$ to $N_{\mathrm{ngb}}=384$. This implies an increase in kernel size by a factor of two. In Fig. 5, we show the results for the mass enclosed within 10 pc and the energy conservation. For almost all of the evolution, there is basically no difference between the two simulations. Only at very late stages, where $h/l$ has increased a great deal at small radii due to the collapse, do we find that the simulation with the larger value for $N_{\mathrm{ngb}}$ reaches high densities earlier. This could be explained by a larger SIDM kernel size leading to an artificially strong effective heat conduction.
We computed- the ratio of kernel size, $h$, to mean free path, $l$, following Fischer et al. (2024a):
@@
Here, $\sigma_{\mathrm{eff}}$ is defined as the effective cross-section given by Eq. 11. In the case of a velocity-independent isotropic cross-section, $\sigma_{\mathrm{eff}}=\sigma$. The result is given for various simulations and times as a function of radius in Fig. 6. We find that a kernel size that is even a few multiples larger than the mean free path appears to be unproblematic (compare also Fig. 3). In contrast, ratios exceeding an order of magnitude appear to be problematic. However, in general, it depends on the simulation set-up how problematic a specific value for $h/l$ is; in particular, it should depend on the velocity dispersion gradient. In this sense, stages of the gravothermal collapse later than those that we have simulated could be even more problematic. Furthermore, we want to add that there might be a mild dependence on how the geometric factor $\Lambda$ for the interaction probability and the drag force is computed (see eqs. 9, 13 and B4 by Fischer et al., 2021a); for example, the chosen kernel function may matter.
Finally, we note that the error arising from a too large value for $h/l$ does not show up in the energy conservation of the simulation. Hence, this is a good example of where energy conservation alone is not enough to assess the simulation’s accuracy.
@@ 3.6 Using a minimal time step
Figure 7: Employing a minimum time step. As in Fig. 1, we show the enclosed mass within 10 pc and the energy conservation for various simulations. This time, we investigate how much the simulations are affected when employing a minimum time step. The brighter lines indicate simulations without a minimum time step, i.e. the time step can decrease further. In contrast the darker lines feature a minimum time step of $\Delta t_{\mathrm{min}}=2.8\times 10^{-5}\,\mathrm{Gyr}$. The displayed simulations J, N, O, Jt, Nt, and Ot are described in Table 1.
@@
Figure 8: Density, velocity dispersion, and Knudsen number for simulations employing a minimum time step. The purple lines give the results for the simulation with $N=2\times 10^{6}$ and $\epsilon=6.4\,\mathrm{pc}$ (grey) and $\Delta t_{\mathrm{min}}=2.8\times 10^{-5}\,\mathrm{Gyr}$ and the blue lines are for the simulation with $N=1\times 10^{7}$ and $\epsilon=1.0\,\mathrm{pc}$ and $\Delta t_{\mathrm{min}}=2.8\times 10^{-5}\,\mathrm{Gyr}$. These are the same simulations as in Fig. 7. The top panel gives the density as a function of radius for different times. We note that these times are during the collapse phase, and for most of them the energy conservation error is sizeable. The middle panel gives the velocity dispersion as a function of radius, and the bottom panel displays the Knudsen number (Eq. (3)). All parameters for the shown simulations Jt and Ot are given in Table 1.
@@
Next, we discuss how imposing a minimum time step impacts the simulation results in the collapse phase. The time step criterion for SIDM, but also gravity, requires smaller and smaller time steps while the halo is collapsing. In particular, the SIDM time step criterion becomes prohibitively small. This poses a challenge for cosmological simulations as one aims to run a simulation until a specific redshift (e.g. $z=0$). A strategy to complete such a simulation, although some halos are collapsing, is to set a minimum allowed time step despite the time step criteria, i.e. evolving some particles on too large time steps.
@@
Following this idea, we set a minimum time step for the late evolution phase. In practice, we continued some of our simulations with this minimum time step. In our case, we chose $\Delta t_{\mathrm{min}}=2.8\times 10^{-5}\,\mathrm{Gyr}$. The results of the simulations for which we tested this are displayed in Fig. 7. Here we show the mass enclosed within 10 pc and the energy conservation error. For the late collapse phase, we find that the simulations with the minimum time step agree well with those that can contain particles on even smaller time steps. Having a minimum time step in place, we were able to continue the simulations to even later stages with higher central densities. However, at some point, the central density stopped growing, and the total energy was no longer conserved, but increased drastically.
Interestingly, the increase in total energy seems to be sensitive to the gravitational softening length and number of particles. For the three runs with the fixed minimum time step, the one with the highest number of particles and the smallest softening length actually has the most severe energy violation. The cause of this surprising behaviour is not entirely clear. Although we impose a minimum time step, the size of the time step at which a particle is evolved can still change. As a consequence, we expect a contribution from the non-time-symmetric change of time steps to the energy conservation error. We note that this error can depend on the choice of the gravitational softening length (see Sect. 3.1). Another source of the energy conservation error comes from the asymmetry of the one-sided oct-tree used in the gravitational force computations (Fischer et al., 2024a). In addition, how quickly the numerical errors grow depends on the size of the employed time step. In particular, the ratio of the actually used time step to the one implied by the time step criteria can matter, and are independent of the adopted minimum time step. Thus, although the minimum time step is fixed to a constant, the time step required by the time step criteria depends effectively on the number of particles and the softening length. The interplay of these factors complicates the evolution of the energy conservation error.
We would like to emphasise that although the runs with relatively low resolution have better energy conservation, it does not mean they are accurate. For example, the simulations shown reach densities for which the SIDM kernel sizes might be relatively large compared to the mean free path, leading to an underestimation of the collapse time, as discussed in Sect. 3.5. To efficiently and accurately simulate such late stages, more work is needed, especially in developing measures to evaluate the accuracy of the simulations, as well as alternative models to describe the inner region of the halo, as we discuss in Sect. 5.
@@
Studying the mass enclosed only within 10 pc may give an incomplete picture of the impact of imposing a minimum time step. Instead, we consider the density, the velocity dispersion and the Knudsen number,
$Kn=\sqrt{\frac{4\uppi}{\rho\,\nu^{2}}}\left(\frac{\sigma_{\mathrm{eff}}}{m}\right)^{-1}\,.$ (3)
@@
We use the effective cross-section, $\sigma_{\mathrm{eff}}$, given by Eq. 11. Moreover, for a velocity-independent isotropic cross-section, $\sigma_{\mathrm{eff}}=\sigma$. We computed those quantities as a function of radius and show them for two of our continued simulations in Fig. 8. For the first time that we show ($4.7\,\mathrm{Gyr}$), the two simulations give fairly similar results. However, at later times they deviate from each other, sometimes leading to huge differences (e.g. for the velocity dispersion). This demonstrates that measuring quantities from simulations of collapsing halos where a minimum time step had been imposed can be far off.
@@
Nevertheless, the simulation Jt shown in purple in Fig. 8 might be more accurate as the energy conservation is much better. The results for the very late collapse phase, where we find an increase in the central density compared to the maximum core formation by a factor of $6\times 10^{5}$, also appear to be qualitatively in agreement with the findings in Fig. 3 by Balberg et al. (2002) for the gravothermal fluid model.
@@
In Sect. 3.7, we comment further on the errors arising from imposing a minimum time step. In detail, we look at the enclosed mass in projection, which is relevant for gravitational lensing.